Operations Report – March 14th

Crew 223 Operations Report 14-MAR-2020

SOL: n/a

Name of person filing report: Shannon Rupert

Non-nominal systems: Nothing to report

Notes on non-nominal systems: Nothing to report

Friday generator:

Off at 6:00am

On at 10:00pm SOC was 86 %

Saturday generator

Off at 6 am

At 7 pm, SOC was 99%

I do not plan to turn it on tonight

Diesel Reading – 60%

Station Propane Reading – 69 %

Ethanol Free Gasoline – 5.5 gallons

Water (loft tank): 40 gallons

Water Meter: n/a units

Water (static tank) – 300 gallons

Static to Loft Pump used – not known

Water in GreenHab – did not check gallons

Water in ScienceDome: 0 gallons

Toilet tank emptied: yes

Perseverance rover used: no

Hours: 152.5

Beginning charge: 100%

Ending charge: 100%

Currently charging: yes

Sojourner rover used: no

Hours: 104.5

Beginning charge: 100%

Ending charge: 100%

Currently charging: yes

Spirit rover: no

Hours: 134.2

Beginning charge: 100%

Ending charge: 100%

Currently charging: yes

Opportunity rover used: no

Hours: 70.3

Beginning charge: 100%

Ending charge: 100%

Currently charging: yes

Curiosity rover used: no

Hours: 139.6

Beginning charge: 100%

Ending charge: 100%

Currently charging: yes

Notes on rovers: Nothing to report

ATV’s Used: No

Reason for use: n/a

Oil Added? no

ATV Fuel Used: None

Hours the ATVs were Used today: 0

Notes on ATVs: Nothing to report.

HabCar used and why, where? Not used.

CrewCar used and why, where? The CrewCar is in Grand Junction and the

crew did not use it at all. They used the director’s truck to go to

Grand Junction this morning and pick up two rental vehicles to drive

to LAX.

General notes and comments: Nothing to report

Summary of internet: All ok but slow due to high winds

Summary of suits and radios: nothing to report

Summary of Hab operations: I turned the heat off at the Hab to conserve power.

Summary of GreenHab operations: I turned the heater down to 16 degrees

to conserve power. I also unplugged the supplemental lights as they

are really not needed anymore and also to conserve power.

Summary of ScienceDome operations: Technician came to see what was

needed to complete installation of dual split unit. This will be

completed next weekend.

Summary of RAM operations: Nothing to report.

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: Nothing to report

Questions, concerns and requests to Mission Support: Happy Pi Day!